Question 14.

You are trying to create an Optimal Complete Portfolio for your client.
To do this you need to estimate the clients Risk Aversion.
You ask a number of questions, comparing different investment alternatives and conclude that the following two alternatives are equally attractive for the client

E(r)

Stdev

Alt 1

12%

20%

Alt 2

5%

7%

a) what is the clients risk aversion?
Answer #1

{\displaystyle A={\frac {dE(c)}{d\sigma }}}

A=RISK AVERSION

dE(c) = change in expected return

d(row) = change in standard deviation

A=(0.12-.005)/(0.20-0.07)

=0.07/0.13

=0.54